What does it mean to refer to a 20 inch Tv set or a 25 inch TV set? Such units refer to the diagonal of the screen. A 10 inch tv set also has a width of 8 inches. What is its height?
Answer by stanbon(75887)

What does it mean to refer to a 20 inch Tv set or a 25 inch TV set? Such units refer to the diagonal of the screen. A 10 inch tv set also has a width of 8 inches. What is its height?
---------------------------------------
Use Pythagoras to findthe height:
height^2 + 8^2 = 10^2
h^2 = 100 - 64
h^2 = 36
h = 6 inches (height)
===========================
Cheers,
Stan H.